the poh of a solution is 4,the H^+ ion concentration will be? 

Grade:12th pass

1 Answers

Vishwas
6 years ago
The h+ conc. Can be found as pH=14-4=10(h+) =10^-10This can be found directly as (OH-) =10^-4And (h+) (oh-) =10^-10